A 6% acidic solution is combined with a 20% acidic solution to produce 21 gallons of a 12% acidic solution. Determine how many gallons of the 6% acidic solution should be used.
Answer by richwmiller(17219)

We need a total of 21 gal at 12%
We want to know how many gal at 6% to mix with 20%
a+b=21,
0.06*a+0.2*b=0.12*21
a=21-b
0.06*(21-b)+0.2*b=2.52
1.26-0.06b+0.2*b=2.52
0.14*b=1.26
b=9
a=21-b
a=12 gal at 6%
b=9 gal at 20%
check
0.06*12+0.2*9=0.12*21
0.72+1.8=2.52
2.52=2.52
ok
codemx1